# Deep Generative Models in Condition and Structural Health Monitoring: Opportunities, Limitations and Future Outlook

### v2

@@ 1.1 Background
Safety and reliability are of great importance for industrial operational systems, driving the need for predictive strategies to mitigate risks and optimize performance. Predictive maintenance (PdM) has emerged as a pivotal strategy, leveraging intelligent data analysis and monitoring techniques to prevent potential failures, minimize downtime, and prolong equipment lifespans [1, 2, 3]. As a core component of PdM, condition monitoring (CM), also termed condition-based maintenance [4], is a real-time process involving sensor-based technologies to track equipment parameters such as vibration, temperature and acoustics in order to assess system health and detect early anomalies. Therefore, CM also involves fault diagnosis and anomaly detection (FD&AD) tasks using real-time data and enables immediate corrective actions to prevent failures. Specifically, in rotating machinery systems, CM stands out by identifying subtle deviations from normal operating profiles, detecting incipient faults (e.g., bearing wear, gear degradation, misalignment due to improper installation, asymmetric loads). These insights are pivotal for rotating systems, where unexpected failures can lead to safety hazards, production losses, or environmental impacts. For static structures, such as pipelines, bridges, or fixed machinery, structural health monitoring (SHM) and non-destructive testing (NDT) are frequently used to assess structural integrity without disrupting operations [5, 6]. NDT techniques, including ultrasonic testing and thermography, assess structural safety without compromising functionality by identifying defects such as cracks or corrosion. SHM expands on this by deploying permanent sensor networks to continuously monitor static systems for gradual degradation.
@@
For intelligent CM/SHM systems, deep neural networks (DNNs) are widely utilized for FD&AD tasks. Representative frameworks, such as convolutional neural networks [7], recurrent neural networks [8, 9], and more recent Transformer-based networks [10], have demonstrated effectiveness in FD&AD tasks by extracting useful features from heterogeneous data modalities, including images, acoustic signals, and time series sensor data. However, limitations and challenges still exist for applying DNNs within industrial scenarios:
@@
Scarce or missing data: Prevalent DNNs are limited by their dependence on large, high-quality labeled datasets to achieve robust classification or prediction performance. Industrial environments frequently struggle to collect sufficient labeled data, or the presence of missing data due to sensor failure, communication interruption, and environmental interference [11, 12];
@@
Domain mismatch: Sensor data collected under normal and abnormal conditions often exhibits severe class imbalance, as industrial systems predominantly operate within nominal states during routine cycles. Even when fault data are available, they may display uneven distribution across distinct fault categories, compromising model training and generalization [13].
@@
Multimodal data complexity: Industrial sensor data originate from heterogeneous sources, such as vibration time series, thermal imaging, acoustic signals, and emission sensors. These multimodal datasets require advanced techniques to address challenges in feature alignment, temporal synchronization, and cross-modal fusion [12].
@@ 1.2 Deep generative models
Deep generative models (DGMs) are deep neural networks with multiple hidden layers trained to approximate complicated, high-dimensional probability distributions of training samples [16, 17]. These models have become the dominant paradigm in GenAI, driven by advances in stochastic optimization and the availability of massive training datasets [18, 19, 20]. The term deep indicates the use of deep neural networks as parametric models in the approximation process [21]. To anchor the view in industrial health monitoring, this review focuses on five prominent DGM frameworks: deep autoregressive (DAR) models, variational autoencoders (VAEs), generative adversarial networks (GANs), diffusion-based models, and large language models (LLMs). Other DGM families, such as normalizing flow models [22], energy-based models [23], and score-based models [24], often integrate with diffusion-based or LLM-based models and are more prevalent in the computer science community. To maintain consistency and avoid misalignment, this review will focus on the aforementioned models, which have already shown promising applications in industrial CM and SHM systems.
Comprehensive scientometric analysis was performed among multiple databases, including Web of Science, Scopus, IEEE Xplore and Google Scholar, to identify relevant publications up to July 2025. Considering that GenAI represents an emerging and rapidly developing research hotspot, numerous studies are available as preprints before being sent to formal peer review. Therefore, the search scope was extended to include repositories such as arXiv, SSRN to capture the most recent trend. The search strategy combined terms related to generative modeling and industrial applications, using keywords such as deep generative model, generative adversarial network, condition monitoring, and structural health monitoring. To ensure completeness, variations and synonyms of these terms (e.g., fault diagnosis, predictive maintenance, and data-driven modeling) were also incorporated. In this way, we provided a broad overview of the evolving scope of generative models applied to industrial health monitoring tasks and systems.
To further elaborate on the co-occurrence among the keywords, we create a map of the cluster of keywords in Fig. 1 using VOSviewer developed by the Center for Science and Technology Studies at Leiden University [25]. The visualization reveals that the generative model exhibits a strong semantic relation with terms such as condition monitoring, fault diagnosis, and data augmentation for vibration signals. However, recent trends indicate a shift toward applications in image processing, synthetic data generation, conditional GANs, and etc. This trend suggests that the deployment of DGMs in industrial CM/SHM remains in its nascent stage, with promising opportunities for future industrial scenarios.
@@ 1.3 Motivations
The remainder of this paper is organized as follows: Section 2 presents the theoretical background and recent state-of-the-art DGMs. Section 3 reviews the tasks and specific industrial applications. Section 4 discusses existing challenges and future research directions. Section 5 concludes the review.
@@ 2.1 Theoretical preliminaries
Mathematically, the primary goal of DGMs involves learning representations of intractable probability distributions over , where is typically large and the distribution exhibits complex characteristics. This learning process uses a finite but substantial collection of independent and identically distributed (i.i.d.) samples from , referred to as training data. Unlike conventional statistical inference that seeks explicit probability expressions, the main objective of DGMs is to construct a generator [16]:
@@ 2.2.1 Deep autoregressive models
Deep autoregressive models belong to a class of DGMs that generate new data samples by iteratively invoking the model based on past observations [26, 27]. This makes them particularly suitable for sequential data modeling tasks such as time series forecasting. Given a dataset with each , the training objective for a DAR model minimizes the negative log-likelihood (NLL) of the data under the model, computed via the chain rule of probability [17]:
@@ 2.2.2 Variational autoencoders
VAEs represent a typical class of DGMs specializing in data compression and reconstruction [17]. Their architecture consists of two networks: an encoder and a decoder. The encoder maps high-dimensional inputs to a lower-dimensional latent space that captures essential data features, while the decoder reconstructs approximations of the original input from these latent representations. This framework achieves effective dimensionality reduction by identifying latent variables that encode the intrinsic structure of data. Formally, a VAE is a probabilistic model that processes an input vector by introducing a latent variable with a predefined prior distribution (e.g., standard Gaussian distribution). Given a dataset with each , the model optimizes the log-likelihood of the data under the generative process:
To tractably maximize this log-likelihood, VAEs optimize the Evidence Lower Bound (ELBO) for each data point , defined as [28]:
@@ 2.2.3 Generative adversarial networks
GANs [29] are a class of DGMs that employ a unique adversarial training process. A GAN consists of two competing networks: a generator (denoted ) and a discriminator (denoted ). The generator tries to create synthetic samples (e.g., thermal images, vibration acceleration signals) given a random noise, while the discriminator attempts to distinguish between whether the generated samples are real or fake [30]. This adversarial process leads the generator to progressively refine its output samples, ultimately producing realistic high-fidelity outputs. The objective function governing this process is formulated as a minimax optimization problem [31]:
@@ 2.2.4 Diffusion models
Denoising diffusion probabilistic models (DDPMs) [32], commonly referred to as diffusion models, form the foundation of the diffusion-based GenAI framework. These models represent a class of DGMs that have demonstrated exceptional performance in image generation tasks. The core of the DDPMs is to construct two Markov chains: a forward diffusion process and its corresponding backward denoising. The forward diffusion process incrementally corrupts an initial data sample over discrete steps by introducing additive Gaussian noise, until it is transformed into a pure noise sample . Reversely, the backward denoising, as it indicates, iteratively reverses this transformation to reconstruct the original data distribution from pure noise. Formally, the forward process defines a Markov chain governed by the forward transition probability [33, 32]:
where denotes the time-dependent noise scale at step , is the identity matrix, and represents the Gaussian distribution. The backward denoising denoted as iteratively reconstructs the original data distribution from pure noise . Specifically, the backward process defines another Markov chain governed by [32]:
where the mean and covariance are learned by a neural network (typically using a U-Net architecture), parameterized by . The loss function of training is to maximize the likelihood by maximizing the ELBO, . The ELBO can be written as matching the true denoising distribution with the parameterized denoising model by [34]:
@@ 2.2.5 Large language models
LLMs have become a dominant paradigm for addressing various engineering challenges. These models are particularly notable for their ability to process diverse inputs, such as text prompts, questions, or visual images through prompt engineering. Representative LLMs such as the generative pre-trained transformer 4 (GPT-4) 111https://openai.com/index/gpt-4/, Llama 4 222https://www.llama.com/, Claude 333https://claude.ai/, and DeepSeek 444https://www.deepseek.com/, all utilize a Transformer-based architecture pre-trained on massive datasets. This architecture integrates attention mechanisms to capture both local and global dependencies between tokens [35]. Mathematically, LLMs function as autoregressive models that predict the next token conditioned on previous tokens, which can be expressed as the following conditional probability:
@@
Specialize for time series or flow data [36]
@@
Struggle to model complex, non-autoregressive distributions [37]
@@
Probabilistic encoder-decoder architecture [38]
@@
May suffer from posterior collapse [39]
@@
Produces high resolution synthetic samples [33]
Prone to training instability [19]
@@
Forward diffusion and backward denoising process [32]
@@
High-fidelity sample generation [26]
Large number of denoising steps to cause training complexity [34]
@@
Multimodal fusion via attention mechanisms [35]
@@
∗ The thermal images of the motor illustrated in the table were retrieved from [41].
@@ 2.3 DGMs vs Classical generative models
Section 2.2 outlines the basic theory of state-of-the-art DGMs. Although generative models have developed rapidly in recent years, some classic methods remain useful and conceptually significant for CM/SHM. Representative classical generative methods include Gaussian mixture models (GMMs) [42], hidden Markov models (HMMs) [43], and autoregressive time-series models such as ARIMA. These are often referred to as statistical or classical generative models. Classical generative models are parametrically compact: a small number of parameters, low computational cost, strong interpretability, and good performance on small datasets as they can explicitly encode domain-specific structures such as linear dynamics or known statistical dependencies. However, their representational capability is limited, as they are often too shallow to capture the highly nonlinear, high-dimensional, or multimodal patterns present in multi-sensor data.
It is worth noting that the autoregressive models span a wide range of models from these classical autoregressive models through deep autoregressive networks to the latest GPT-style autoregressive large models (i.e., LLMs) [44]. On that basis, we compare classical generative models and DGMs with respect to model capacity, data and computational requirements, inference speed, and interpretability, as shown in Table 2.
Metric Classical generative models Deep generative models (GMM, HMM, ARIMA, etc.) (DAR, VAE, GAN, Diffusion, LLMs, etc.) Training objective Maximum likelihood, Expectation–Maximization, least squares ELBO (VAE), adversarial loss(GAN), denoising score (Diffusion model) Data types Low dimension data, time series High-dimensional complex data: images, audio, multimodal sensor data Model complexity Typical range: [45] Typical range: (VAEs/GANs [46], transformers/LLMs [47]) Evaluation metrics Log-likelihood, RMSE, prediction interval coverage NLL, ELBO, Fréchet inception distance, task-specific scores (F1-Score, AUC, etc.) Inference speed , typically on CPU [48] Normally on GPU; Diffusion/LLM inference: [49] Interpretability Highly interpretable Low/moderate interpretability due to a large amount of model parameters
@@ 3.1.2 DGMs for data generation
Data generation serves as an umbrella concept with data augmentation, reconstruction and imputation, which differ mainly in how and why the data are generated. Data augmentation can improve model robustness by generating new samples from existing data, while data reconstruction tries to restore or rebuild signal from partial or compressed data. In contrast, data imputation focuses on filling in missing or corrupted values to complete the dataset. Following the data augmentation methods defined in [50], these techniques can be categorized into three primary groups: the synthetic minority over-sampling technique (SMOTE), data transformation-based augmentation, and generative model-based approaches. SMOTE-based methods specifically address class imbalance by synthetically generating minority class samples (e.g., fault or defect samples in CM/SHM applications). Whereas data transformation-based approaches enhance dataset diversity through operations such as noise injection or image manipulations (e.g., cropping, rotating, flipping, and padding). Fig. 4 is an illustration for both 1D time-series and 2D image augmentation in CM/SHM applications. The third category comprises generative model-based approaches, which in this work refer to DGM-based approaches.
@@
DGM-based data augmentation methods have emerged as indispensable tools to address challenges in scenarios where real-world data is scarce and noisy. These models enable robust supplementation and augmentation of existing datasets by synthesizing high-quality data that mirrors the statistical properties of real data samples, especially when acquiring experimental operation data is often prohibitively costly and labor-intensive [53, 54]. Thus, by adopting DGM-based data augmentation, industrial sectors are expected to mitigate data scarcity and address imbalance issues inherent to industrial settings, thus enriching training datasets for downstream FD&AD models while preserving the integrity of original data distributions [55, 56]. Beyond data augmentation, DGMs also play a critical role in data imputation, a solution to resolve missing data points in industrial systems [12]. Missing data may seriously compromise the diagnostic or detective performance of DGM-based models [57]. The inherent ability of DGMs to learn complex data distributions makes them well-suited for data imputation tasks. Instead of simply filling in missing values with fixed estimates, DGMs can leverage the underlying patterns and correlations present in the data to generate plausible values for missing entries. Table 4 summarizes the existing literature on applying DGMs to data augmentation and imputation tasks. We comprehensively evaluate signal processing, conventional ML-based approaches (e.g., support vector regression (SVR), recurrent neural networks (RNN), long short-term memory (LSTM), gated recurrent unit (GRU)), and DGM-based methods using the performance metrics outlined in Table 4. SMOTE-based approaches are classified as conventional ML, as SMOTE generates synthetic data through linear interpolation between minority class samples and their k-nearest neighbors (KNN) [58].
@@
Existing studies explicitly focused on signal processing, conventional ML, and DGM are listed in Refs. Works comparing the performance improvements across these approaches using evaluation metrics in Table 4 are detailed in Quantitative comparison. From the table, it is found: (i) most works focus exclusively on data augmentation or imputation without cross-comparative analysis; (ii) in quantitative comparisons, DDPM and GAN-based models consistently outperform conventional ML methods (e.g., LSTM, SMOTE) across metrics such as MMD and EMD; (iii) cross comparisons between signal processing methods, conventional DNNs, and DGMs remain scarce, largely because signal processing techniques are primarily employed as preprocessing stages.
@@
Moreover, diffusion-based models are typically large-scale architectures containing billions of parameters. Recent studies have examined the relationship between the training time and data generation quality in diffusion-based models [59, 60, 61]. As discussed in these works, the number of diffusion steps is a key factor determining the final data generation quality, which typically ranges from 500 to 1000 steps [61]. This observation has motivated the development of denoising diffusion implicit models (DDIMs) [34], which aim to accelerate the sampling process while maintaining generation fidelity. In practice, the training time for a typical diffusion model often exceeds 10 hours even when using commercial-grade GPUs.

### v3

@@ 1.1 Background
Safety and reliability are of great importance for industrial operational systems, driving the need for predictive strategies to mitigate risks and optimize performance. Predictive maintenance (PdM) has emerged as a pivotal strategy, leveraging intelligent data analysis and monitoring techniques to prevent potential failures, minimize downtime, and prolong equipment lifespans [NUNES202353, 10440027, huxiaosong2021]. As a core component of PdM, condition monitoring (CM), also termed condition-based maintenance [AHMEDMURTAZA2024102935], is a real-time process involving sensor-based technologies to track equipment parameters such as vibration, temperature and acoustics in order to assess system health and detect early anomalies. Therefore, CM also involves fault diagnosis and anomaly detection (FD&AD) tasks using real-time data and enables immediate corrective actions to prevent failures. Specifically, in rotating machinery systems, CM stands out by identifying subtle deviations from normal operating profiles, detecting incipient faults (e.g., bearing wear, gear degradation, misalignment due to improper installation, asymmetric loads). These insights are pivotal for rotating systems, where unexpected failures can lead to safety hazards, production losses, or environmental impacts. For static structures, such as pipelines, bridges, or fixed machinery, structural health monitoring (SHM) and non-destructive testing (NDT) are frequently used to assess structural integrity without disrupting operations [YANG2024111597, AliMardanshahi2025]. NDT techniques, including ultrasonic testing and thermography, assess structural safety without compromising functionality by identifying defects such as cracks or corrosion. SHM expands on this by deploying permanent sensor networks to continuously monitor static systems for gradual degradation.
@@
For intelligent CM/SHM systems, deep neural networks (DNNs) are widely utilized for FD&AD tasks. Representative frameworks, such as convolutional neural networks [CHEN2019106272], recurrent neural networks [CHEN2020106683, 2022_Yang], and more recent Transformer-based networks [WU2023439], have demonstrated effectiveness in FD&AD tasks by extracting useful features from heterogeneous data modalities, including images, acoustic signals, and time series sensor data. However, limitations and challenges still exist for applying DNNs within industrial scenarios:
@@
Scarce or missing data: Prevalent DNNs are limited by their dependence on large, high-quality labeled datasets to achieve robust classification or prediction performance. Industrial environments frequently struggle to collect sufficient labeled data, or the presence of missing data due to sensor failure, communication interruption, and environmental interference [VANDREVEN2024132711, ZHANG2025110663];
@@
Domain mismatch: Sensor data collected under normal and abnormal conditions often exhibits severe class imbalance, as industrial systems predominantly operate within nominal states during routine cycles. Even when fault data are available, they may display uneven distribution across distinct fault categories, compromising model training and generalization [ZHAO2024122807].
@@
Multimodal data complexity: Industrial sensor data originate from heterogeneous sources, such as vibration time series, thermal imaging, acoustic signals, and emission sensors. These multimodal datasets require advanced techniques to address challenges in feature alignment, temporal synchronization, and cross-modal fusion [ZHANG2025110663].
Generalization limitations: Current FD&AD methods often lack adaptability and stability when deployed across varying operational environments, equipment configurations, or operating conditions, restricting their practical applicability in dynamic industrial scenarios [ZHANG2025124378, hu2024evaluationimprovement].
@@ 1.2 Deep generative models
Deep generative models (DGMs) are deep neural networks with multiple hidden layers trained to approximate complicated, high-dimensional probability distributions of training samples [ruthotto2021introduction, Bond_Taylor_2022]. These models have become the dominant paradigm in GenAI, driven by advances in stochastic optimization and the availability of massive training datasets [manduchi2025chall, ZHANG2025125059, GM2020100285]. The term deep indicates the use of deep neural networks as parametric models in the approximation process [Tomczak2022]. To anchor the view in industrial health monitoring, this review focuses on five prominent DGM frameworks: deep autoregressive (DAR) models, variational autoencoders (VAEs), generative adversarial networks (GANs), diffusion-based models, and large language models (LLMs). Other DGM families, such as normalizing flow models [Kobyzev_2021], energy-based models [NEURIPS2019_378a063b], and score-based models [song2021scorebasedgenerative], often integrate with diffusion-based or LLM-based models and are more prevalent in the computer science community. To maintain consistency and avoid misalignment, this review will focus on the aforementioned models, which have already shown promising applications in industrial CM and SHM systems.
Comprehensive scientometric analysis was performed among multiple databases, including Web of Science, Scopus, IEEE Xplore and Google Scholar, to identify relevant publications up to the end of 2025. Considering that GenAI represents an emerging and rapidly developing research hotspot, numerous studies are available as preprints before being sent to formal peer review. Therefore, the search scope was extended to include repositories such as arXiv, SSRN to capture the most recent trend. The search strategy combined terms related to generative modeling and industrial applications, using keywords such as deep generative model, generative adversarial network, condition monitoring, and structural health monitoring. To ensure completeness, variations and synonyms of these terms (e.g., fault diagnosis, predictive maintenance, and data-driven modeling) were also incorporated. In this way, we provided a broad overview of the evolving scope of generative models applied to industrial health monitoring tasks and systems.
To further elaborate on the co-occurrence among the keywords, we create a map of the cluster of keywords in Fig. 1 using VOSviewer developed by the Center for Science and Technology Studies at Leiden University [van2010software]. The visualization reveals that the generative model exhibits a strong semantic relation with terms such as condition monitoring, fault diagnosis, and data augmentation for vibration signals. However, recent trends indicate a shift toward applications in image reconstruction, synthetic data generation, diffusion model, and etc. This trend suggests that the deployment of DGMs in industrial CM/SHM remains in its nascent stage, with promising opportunities for future industrial scenarios.
@@ 1.3 Motivations
The remainder of this paper is organized as follows: Section 2 presents the theoretical background and recent state-of-the-art DGMs. Section 3 reviews the tasks and specific industrial applications. Section LABEL:Sec:4 discusses existing challenges and future research directions. Section LABEL:Sec:5 concludes the review.
@@ 2.1 Theoretical preliminaries
Mathematically, the primary goal of DGMs involves learning representations of intractable probability distributions over , where is typically large and the distribution exhibits complex characteristics. This learning process uses a finite but substantial collection of independent and identically distributed (i.i.d.) samples from , referred to as training data. Unlike conventional statistical inference that seeks explicit probability expressions, the main objective of DGMs is to construct a generator [ruthotto2021introduction]:
@@ 2.2.1 Deep autoregressive models
Deep autoregressive models belong to a class of DGMs that generate new data samples by iteratively invoking the model based on past observations [urain2024dgmsrobotics, ren2025aigcindustrial]. This makes them particularly suitable for sequential data modeling tasks such as time series forecasting. Given a dataset with each , the training objective for a DAR model minimizes the negative log-likelihood (NLL) of the data under the model, computed via the chain rule of probability [Bond_Taylor_2022]:
@@ 2.2.2 Variational autoencoders
VAEs represent a typical class of DGMs specializing in data compression and reconstruction [Bond_Taylor_2022]. Their architecture consists of two networks: an encoder and a decoder. The encoder maps high-dimensional inputs to a lower-dimensional latent space that captures essential data features, while the decoder reconstructs approximations of the original input from these latent representations. This framework achieves effective dimensionality reduction by identifying latent variables that encode the intrinsic structure of data. Formally, a VAE is a probabilistic model that processes an input vector by introducing a latent variable with a predefined prior distribution (e.g., standard Gaussian distribution). Given a dataset with each , the model optimizes the log-likelihood of the data under the generative process:
To tractably maximize this log-likelihood, VAEs optimize the Evidence Lower Bound (ELBO) for each data point , defined as [SHI2023101940]:
@@ 2.2.3 Generative adversarial networks
GANs [NIPS2014_f033ed80] are a class of DGMs that employ a unique adversarial training process. A GAN consists of two competing networks: a generator (denoted ) and a discriminator (denoted ). The generator tries to create synthetic samples (e.g., thermal images, vibration acceleration signals) given a random noise, while the discriminator attempts to distinguish between whether the generated samples are real or fake [wen2024gear]. This adversarial process leads the generator to progressively refine its output samples, ultimately producing realistic high-fidelity outputs. The objective function governing this process is formulated as a minimax optimization problem [he2024llmsmeetmultimodalgeneration]:
@@ 2.2.4 Diffusion models
Denoising diffusion probabilistic models (DDPMs) [ho2020denoisingdiffusion], commonly referred to as diffusion models, form the foundation of the diffusion-based GenAI framework. These models represent a class of DGMs that have demonstrated exceptional performance in image generation tasks. The core of the DDPMs is to construct two Markov chains: a forward diffusion process and its corresponding backward denoising. The forward diffusion process incrementally corrupts an initial data sample over discrete steps by introducing additive Gaussian noise, until it is transformed into a pure noise sample . Reversely, the backward denoising, as it indicates, iteratively reverses this transformation to reconstruct the original data distribution from pure noise. Formally, the forward process defines a Markov chain governed by the forward transition probability [xiao2022tacklinggenerative, ho2020denoisingdiffusion]:
where denotes the time-dependent noise scale at step , is the identity matrix, and represents the Gaussian distribution. The backward denoising denoted as iteratively reconstructs the original data distribution from pure noise . Specifically, the backward process defines another Markov chain governed by [ho2020denoisingdiffusion]:
where the mean and covariance are learned by a neural network (typically using a U-Net architecture), parameterized by . The loss function of training is to maximize the likelihood by maximizing the ELBO, . The ELBO can be written as matching the true denoising distribution with the parameterized denoising model by [song2022ddim]:
@@ 2.2.5 Large language models
LLMs have become a dominant paradigm for addressing various engineering challenges. These models are particularly notable for their ability to process diverse inputs, such as text prompts, questions, or visual images through prompt engineering. Representative LLMs such as the generative pre-trained transformer 4 (GPT-4) 111https://openai.com/index/gpt-4/, Llama 4 222https://www.llama.com/, Claude 333https://claude.ai/, and DeepSeek 444https://www.deepseek.com/, all utilize a Transformer-based architecture pre-trained on massive datasets. This architecture integrates attention mechanisms to capture both local and global dependencies between tokens [vaswani2023attentionneed]. Mathematically, LLMs function as autoregressive models that predict the next token conditioned on previous tokens, which can be expressed as the following conditional probability:
@@
Specialize for time series or flow data [Yanfei2020]
@@
Struggle to model complex, non-autoregressive distributions [ye2025autoregression]
@@
Probabilistic encoder-decoder architecture [Kingma_2019]
@@
May suffer from posterior collapse [Lucas2019UnderstandingPC]
@@
Produces high resolution synthetic samples [xiao2022tacklinggenerative]
Prone to training instability [ZHANG2025125059]
@@
Forward diffusion and backward denoising process [ho2020denoisingdiffusion]
@@
High-fidelity sample generation [urain2024dgmsrobotics]
Large number of denoising steps to cause training complexity [song2022ddim]
@@
Multimodal fusion via attention mechanisms [vaswani2023attentionneed]
@@
∗ The thermal images of the motor illustrated in the table were retrieved from [NajafiDataset].
@@ 2.3 DGMs vs Classical generative models
Section 2.2 outlines the basic theory of state-of-the-art DGMs. Although generative models have developed rapidly in recent years, some classic methods remain useful and conceptually significant for CM/SHM. Representative classical generative methods include Gaussian mixture models (GMMs) [WEN2024107562], hidden Markov models (HMMs) [CORACA2023109025], and autoregressive time-series models such as ARIMA. These are often referred to as statistical or classical generative models. Classical generative models are parametrically compact: a small number of parameters, low computational cost, strong interpretability, and good performance on small datasets as they can explicitly encode domain-specific structures such as linear dynamics or known statistical dependencies. However, their representational capability is limited, as they are often too shallow to capture the highly nonlinear, high-dimensional, or multimodal patterns present in multi-sensor data.
It is worth noting that the autoregressive models span a wide range of models from these classical autoregressive models through deep autoregressive networks to the latest GPT-style autoregressive large models (i.e., LLMs) [naveed2024comprehensiveoverviewlargelanguage]. On that basis, we compare classical generative models and DGMs with respect to model capacity, data and computational requirements, inference speed, and interpretability, as shown in Table 2.
Metric Classical generative models Deep generative models (GMM, HMM, ARIMA, etc.) (DAR, VAE, GAN, Diffusion, LLMs, etc.) Training objective Maximum likelihood, Expectation–Maximization, least squares ELBO (VAE), adversarial loss(GAN), denoising score (Diffusion model) Data types Low dimension data, time series High-dimensional complex data: images, audio, multimodal sensor data Model complexity Typical range: [3408318] Typical range: (VAEs/GANs [hibat2024framework], transformers/LLMs [zhang2024scalingmeetsllmfinetuning]) Evaluation metrics Log-likelihood, RMSE, prediction interval coverage NLL, ELBO, Fréchet inception distance, task-specific scores (F1-Score, AUC, etc.) Inference speed , typically on CPU [Zhu_2022] Normally on GPU; Diffusion/LLM inference: [NEURIPS2024_9ad996b5] Interpretability Highly interpretable Low/moderate interpretability due to a large amount of model parameters
@@ 3.1.2 DGMs for data generation
Data generation serves as an umbrella concept with data augmentation, reconstruction and imputation, which differ mainly in how and why the data are generated. Data augmentation can improve model robustness by generating new samples from existing data, while data reconstruction tries to restore or rebuild signals from partial or compressed data. In contrast, data imputation focuses on filling in missing or corrupted values to complete the dataset. Following the data augmentation methods defined in [Ju_2024], these techniques can be categorized into three primary groups: the synthetic minority over-sampling technique (SMOTE), data transformation-based augmentation, and generative model-based approaches. SMOTE-based methods specifically address class imbalance by synthetically generating minority class samples (e.g., fault or defect samples in CM/SHM applications). Whereas data transformation-based approaches enhance dataset diversity through operations such as noise injection or image manipulations (e.g., cropping, rotating, flipping, and padding). Fig. 4 is an illustration for both 1D time-series and 2D image augmentation in CM/SHM applications. The third category comprises generative model-based approaches, which in this work refer to DGM-based approaches.
@@
DGM-based data augmentation methods have emerged as indispensable tools to address challenges in scenarios where real-world data is scarce and noisy. These models enable robust supplementation and augmentation of existing datasets by synthesizing high-quality data that mirrors the statistical properties of real data samples, especially when acquiring experimental operation data is often prohibitively costly and labor-intensive [CHEN2024128167, Zhao3649447]. Thus, by adopting DGM-based data augmentation, industrial sectors are expected to mitigate data scarcity and address imbalance issues inherent to industrial settings, thus enriching training datasets for downstream FD&AD models while preserving the integrity of original data distributions [LIU2024124511, LIU2021107488]. Beyond data augmentation, DGMs also play a critical role in data imputation, a solution to resolve missing data points in industrial systems [ZHANG2025110663]. Missing data may seriously compromise the diagnostic or detective performance of DGM-based models [ma2021identifiable]. The inherent ability of DGMs to learn complex data distributions makes them well-suited for data imputation tasks. Instead of simply filling in missing values with fixed estimates, DGMs can leverage the underlying patterns and correlations present in the data to generate plausible values for missing entries. Table 4 summarizes the existing literature on applying DGMs to data augmentation and imputation tasks. We comprehensively evaluate signal processing, conventional ML-based approaches (e.g., support vector regression (SVR), recurrent neural networks (RNN), long short-term memory (LSTM), gated recurrent unit (GRU)), and DGM-based methods using the performance metrics outlined in Table 4. SMOTE-based approaches are classified as conventional ML, as SMOTE generates synthetic data through linear interpolation between minority class samples and their k-nearest neighbors (KNN) [elreedy2024theoretical].
@@
Existing studies explicitly focused on signal processing, conventional ML, and DGM are listed in Refs. Works comparing the performance improvements across these approaches using evaluation metrics in Table 4 are detailed in Quantitative comparison. From the table, it is found: (i) most works focus exclusively on data augmentation or imputation without cross-comparative analysis; (ii) in quantitative comparisons, DDPM and GAN-based models consistently outperform conventional ML methods (e.g., LSTM, SMOTE) across metrics such as MMD and EMD; (iii) cross comparisons between signal processing methods, conventional DNNs, and DGMs remain scarce. This is largely because signal processing techniques are typically used as preprocessing steps, while DGMs are generally regarded as complementary tools for data generation that do not explicitly incorporate physical knowledge.
@@
Moreover, diffusion-based models are typically large-scale architectures containing billions of parameters. Recent studies have examined the relationship between the training time and data generation quality in diffusion-based models [HU2025113122, SHU2025111783, YANG2025112996]. As discussed in these works, the number of diffusion steps is a key factor determining the final data generation quality, which typically ranges from 500 to 1000 steps [YANG2025112996]. This observation has motivated the development of denoising diffusion implicit models (DDIMs) [song2022ddim], which aim to accelerate the sampling process while maintaining generation fidelity. In practice, the training time for a typical diffusion model often exceeds 10 hours even when using commercial-grade GPUs.